# Flyby-induced displacement effect: an analytic solution

### v4

@@ Abstract
We describe analytically the scattering of particles by a sandwich gravitational wave generated during a flyby. The derivative-of-the-Gaussian profile proposed by Gibbons and Hawking is approximated by the hyperbolic Scarf potential, which allows for an exact analytic solution via the Nikiforov-Uvarov method. Our results confirm the prediction of Zel’dovich and Polnarev when the wave amplitude takes certain “magical” values.
@@ I Introduction
One of the early proposals to observe gravitational waves , put forward by Braginsky and Thorne BraTho , was to study the motion of particles initially in rest after the passing of a burst of gravitational waves . They argued indeed that Burst With Memory could be among the earliest kinds of gravitational waves detected BraTho .
Initial studies Ehlers considered the Velocity Memory (VM) effect : after the wave has passed, the particles would fly apart with constant velocity along straight trajectories. Zel’dovich and Polnarev ZelPol advocated instead in favor of flyby: gravitational waves generated by a cluster of superdense stars (neutron stars or black holes) would yield the Displacement Memory (DM) effect : “…although the distance between a pair of two noninteracting bodies (such as satellites) will change, their relative velocity will become vanishingly small as the flyby concludes.”
The GW memory effect offers a way to test general relativity including the nonlinear nature of gravity, and provides additional information about the GW source. Preliminary studies have suggested that the detection of GW memory from individual mergers of massive binary black holes is expected with LISA Favata . However, more realistic estimates show that observing the GW memory effect at detectors such as LISA, although feasible, is a challenging task fly-1 ; fly-2 . Recent attempts to detect the effect in LIGO/Virgo’s first transient gravitational-wave catalog, as well as in a pulsar timing array experiment NANOGrav, have predictably yielded negative results fly-3 ; fly-4 . However the prospects for the next generation of gravitational wave detectors are brighter even for future ground-based detectors such as Cosmic Explorer and the Einstein Telescope (see fly-2 ; fly-4 and references therein).
Zel’dovich and Polnarev did not present any proof. Although their suggestion could be confirmed in Christo ; BMS1 ; BMS2 ; BMS3 ; BMS4 , there were some doubts due to the apparent contradiction with the Bondi-Pirani theorem on the caustic property of plane gravitational waves BoPi89 ; ShortMemory . Then it came as a surprise that that the DM effect could be confirmed numerically DM-1 ; Jibril ; DM-2 for plane waves — provided the amplitude takes some “magical” values. Below we present an analytic proof.
@@
$ds^{2}=d{\bm{X}}^{2}+2dUdV+{\cal A}(U)\big{(}X_{+}^{2}-X_{-}^{2}\big{)}dU^{2}\,.$ (1)
Gibbons and Hawking GibbHaw71 argued that in linear theory the Riemann tensor of a gravitational wave is proportional to the fourth time derivative of the quadrupole moment of the source GibbonsP . This lead them to propose the derivative of the Gaussian,
@@
as profile. The geodesic equations for (2) do not support any analytic solution, the profile can however be approximated by the derived Pöschl-Teller (dPT) potential PTeller ; Chakra ; DM-1 ,
@@
In this note we present a third approximation of the flyby profile (2), namely by the hyperbolic Scarf potential Scarfpot  [^4] The trigonometric Scarf potential was first introduced in Scarf to mimic the potential of a real crystal. It turned out to be the tip of the iceberg of a whole class of potentials with interesting properties, ranging from classical and quantum integrability Perelomov ; Perelomov1 to supersymmetric quantum mechanics Gendenshtein ; Natanson ; Dutt86 ; Levai ; Scrf1 . ,
@@
to be compared with (3). For small values of $U$ the Scarf profile (4) behaves as the derivative of the Gaussian or of Pöschl-Teller, (2) or (3), but for large $U$ it falls off more slowly. However dilating $U$ appropriately almost perfect overlapping is obtained, as depicted in FIG.1. Their similarity follows from that all three of them are derivatives of similar bell-shaped functions. We note also all three profiles in (2)-(3)-(4) are derivatives of an even function and are therefore odd.
Figure 1: The derived Pöschl-Teller  profile (3) and its Scarf potential alter-ego (4) are both good approximations of the derived-Gaussian one proposed by Gibbons and Hawking GibbHaw71 to describe flyby.
@@
$\dfrac{d^{2}\!X^{\pm}}{dU^{2}}\mp\frac{1}{2}{\cal A}X^{\pm}=0\,,$ (5)
where the $\pm$ label refers to the upper resp. lower component. The sign of the second  term comes from the relative minus between the coordinates in (1). The equation for the $V$ component follows from these DM-1 ; DM-2 , and is therefore omitted. We recall that $U$ behaves, in the Bargmann framework DBKP ; DGH91 , as non-relativistic time. These Sturm-Liouville-type equations are completed with the DM boundary conditions,
@@
Our main result then says that for the profile (4) analytic solutions can be found which confirm the prediction of Zel’dovich and Polnarev ZelPol , supplemented by the quantization condition mentioned above. Our clue is that the geodesic equation (5) can be viewed as a time-independent Schrödinger equation and the DM trajectories correspond to its real (non-normalizable) zero-energy ground states DM-1 .
@@ II Analytic solution for flyby
Finding the geodesics for dPT is considerably more difficult, though, as for “underived” Pöschl-Teller  ; it can be obtained in terms of confluent Heun functions DM-1 ; DM-2 with numerically calculated approximate expressions, see eqn. # (IV.16) of DM-2 .
@@
$\;\;\displaystyle{X_{n}(t)=(1+t^{2})^{-\frac{n}{2}}\,e^{-\mathrm{sign}(g)\,\sqrt{n(n+1)}\arctan{t}}\,y_{n}(t)\,,}\;\;$ (11)
@@ III Transverse geodesics by the Nikiforov-Uvarov method
$\sigma_{3}=\frac{g^{2}}{4k}\left[t^{2}+\frac{4k}{g}t+\frac{4k^{2}}{g^{2}}\right]=\frac{g^{2}}{4k}\left(t+\frac{2k}{g}\right)^{2}\,,\quad\pi=\frac{t}{2}\pm\sqrt{\sigma_{3}}=\frac{t}{2}\pm\frac{|g|}{2\sqrt{k}}\left|t+\frac{2k}{g}\right|\,.\quad$ (17)
@@
Therefore, the corresponding geodesic trajectory is given by eqn. (11). The polynomial $y_{n}(t)$ is given by the Rodrigues formula Eq. (38) which requires to calculate the weight function $\rho(t)$ in (13). Then we end up with eqn. (12).
@@ IV Conclusion
The profile (2) proposed by Hawking and Gibbons for flyby GibbHaw71 can be approximated by both the derived-Pöschl-Teller (3) and by the Scarf potentials (4), as illustrated in FIG.1. The advantage of Scarf is that it supports analytic solutions, (11)-(12), labeled by an integer $n$, eqn (10). They are constructed by the Nikiforov-Uvarov method.
The displacement effect (DM) arises when the wave zone contains $n$ half-waves as observed made DM-2 for the Gaussian and Pöschl-Teller profiles. The number of half-waves is equal to that of the zeros of the polynomial $y_{n}(t)$.
Our Scarf solution is similar to but slightly different from those found semi-analytically for derived-Pöschl-Teller DM-2 , as seen in FIG.2. It has also the advantage to be exact, while the one in DM-2 is only approximate, with error increasing for higher wave numbers.
@@
A recent paper MSUSY confirms our results using supersymmetric quantum mechanics with this superpotential.

### v5

@@ Abstract
The motion of particles hit by a burst of gravitational waves generated by flyby admits, for the derivative-of-the-Gaussian profile, only a numerical description. The profile can however be approximated by the hyperbolic Scarf potential which admits an exact analytic solution via the Nikiforov-Uvarov method. Our toy model is consistent with the prediction of Zel’dovich and Polnarev provided the wave amplitude takes certain “magical” values.
Phys. Lett. B 868 (2025) 139687.  https://doi.org/10.1016/j.physletb.2025.139687
@@ I Introduction
One of the early proposals to observe gravitational waves , put forward by Braginsky and Thorne BraTho , was to study the motion of particles initially in rest after the passing of a burst of gravitational waves . They argued indeed that Burst With Memory could be among the earliest kinds of gravitational waves detected BraTho . Since then, and with the discovery that gravitational memory effects are closely related to soft graviton theorems and Bondi-Metzner-Sachs symmetries at future null infinity, the field has grown considerably BMS1 ; BMS2 ; BMS3 ; BMS4 .
Initial studies Ehlers ; GWGNature considered the Velocity Memory (VM) effect : after the wave has passed, the particles would fly apart with constant velocity along straight trajectories. Zel’dovich and Polnarev ZelPol advocated instead in favor of flyby: gravitational waves generated by a cluster of superdense stars (neutron stars or black holes) would yield the Displacement Memory (DM) effect : “…although the distance between a pair of two noninteracting bodies (such as satellites) will change, their relative velocity will become vanishingly small as the flyby concludes.”
The GW memory effect offers a way to test general relativity including the nonlinear nature of gravity, and provides additional information about the GW source. Preliminary studies have suggested that the detection of GW memory from individual mergers of massive binary black holes is expected with LISA Favata . See Blanchet for a state-of-the-art review. However, more realistic estimates show that observing the GW memory effect at detectors such as LISA, although feasible, is a challenging task fly-1 ; fly-2 . Recent attempts to detect the effect in LIGO/Virgo’s first transient gravitational-wave catalog, as well as in a pulsar timing array experiment NANOGrav, have, predictably, yielded negative results fly-3 ; fly-4 . However the prospects for the next generation of gravitational wave detectors are brighter even for future ground-based detectors such as Cosmic Explorer and the Einstein Telescope (see fly-2 ; fly-4 and references therein).
Although the displacement effect could be confirmed Christo ; BMS1 ; BMS2 ; BMS3 ; BMS4 , there remained some doubts due to the apparent contradiction with the Bondi-Pirani theorem on the caustic property of plane gravitational waves BoPi89 ; ShortMemory .
More recently, the DM effect could be verified numerically for plane waves using the approach proposed in LongMemory , provided the amplitude takes some “magical” values DM-1 ; Jibril ; DM-2 . In this note we study the Zel’dovich - Polnarev statement by using the simplified framework in ref LongMemory . Although they should be viewed only as toy models, they do provide some insight into the memory effect.
Based Gibbons and Hawking’s earlier work Gibbons_1976 , the model in LongMemory proposes that the Brinkmann profile of the gravitational wave generated by flyby be approximated by the first derivative of the Gaussian (dG), eqn. (2) below. This yields however Sturm-Liouville-type geodesic equations with no analytic solution. In this paper we propose a similar but slightly different approximation, namely by the Scarf potential Scarf ; Scarfpot which does provide us with analytic solutions. The results are in full agreement with ZelPol .
@@
$ds^{2}=d{\bm{X}}^{2}+2dUdV+{\cal A}(U)\big{(}X_{+}^{2}-X_{-}^{2}\big{)}dU^{2}\,,$ (1)
where ${\bm{X}}$ is referred to as transverse coordinate; $U$ and $V$ are lightlike. In the non-relativistic Kaluza-Klein-type framework of ref. DBKP ; DGH91 $U$ plays the role of non-relativistic time. Ref. LongMemory argues that in linear theory the Riemann tensor of a gravitational wave is proportional to the fourth time derivative of the quadrupole moment of the source Gibbons_1976 ; GibbonsP . This lead us to propose the derivative of the Gaussian,
@@
as profile. The geodesic equations for (2) have no analytic solution, though, the profile can however be approximated by the derived Pöschl-Teller (dPT) potential PTeller ; Chakra ; DM-1 ,
@@
In this note we approximate the flyby profile (2) by the hyperbolic Scarf potential Scarfpot  [^4] The trigonometric Scarf potential was initially introduced in Scarf to mimic the potential of a real crystal. It turned out to be the tip of the iceberg of a whole class of potentials with interesting properties, ranging from classical and quantum integrability Perelomov ; Perelomov1 to supersymmetric quantum mechanics Gendenshtein ; Natanson ; Dutt86 ; Levai ; Scrf1 ; MSUSY . ,
@@
to be compared with (3). The difference is merely the exponent in the denominator. For small values of $U$ the Scarf profile (4) behaves as the derivative of the Gaussian or of Pöschl-Teller, (2) or (3), but for large $U$ it falls off more slowly. However dilating $U$ appropriately almost perfect overlapping is obtained, as depicted in FIG.1. Their similarity follows from that all three of them are derivatives of similar bell-shaped functions. We note also all three profiles in (2)-(3)-(4) are derivatives of an even function and are therefore odd.
Figure 1: The derived Pöschl-Teller  profile (3) and its Scarf potential alter-ego (4) are both approximations of the derived-Gaussian one proposed in LongMemory to describe flyby.
@@
$\dfrac{d^{2}\!X^{\pm}}{dU^{2}}\mp\frac{1}{2}{\cal A}(U)X^{\pm}=0\,,$ (5)
where the $\pm$ label refers to the upper resp. lower component. The sign of the second  term comes from the relative minus between the coordinates in (1). These Sturm-Liouville-type equations are completed with the DM boundary conditions,
@@
Our main result then says that for the profile (4) analytic solutions can be found, consistently with the prediction of Zel’dovich and Polnarev ZelPol , supplemented by the quantization condition mentioned above. Our clue is that the geodesic equation (5) can be viewed as a time-independent Schrödinger equation and the DM trajectories correspond to its real (non-normalizable) zero-energy ground states DM-1 .
@@ II Analytic solution for flyby
Finding the geodesics for dPT is considerably more difficult, though, as for “underived” Pöschl-Teller  ; it can be obtained in terms of confluent Heun functions DM-2 with numerically calculated approximate expressions, see eqn. # (IV.16) of DM-2 .
@@
$X_{n}(t)=(1+t^{2})^{-\frac{n}{2}}\,e^{-\mathrm{sign}(g)\,\sqrt{n(n+1)}\arctan{t}}\,y_{n}(t)\,,$ (11)
@@ III Transverse geodesics by the Nikiforov-Uvarov method
$\sigma_{3}=\frac{g^{2}}{4k}\left[t^{2}+\frac{4k}{g}t+\frac{4k^{2}}{g^{2}}\right]=\frac{g^{2}}{4k}\left(t+\frac{2k}{g}\right)^{2},\;\pi=\frac{t}{2}\pm\sqrt{\sigma_{3}}=\frac{t}{2}\pm\frac{|g|}{2\sqrt{k}}\left|t+\frac{2k}{g}\right|\,.\>$ (17)
@@
Therefore, the corresponding geodesic trajectory is given by eqn. (11). The polynomial $y_{n}(t)$ is given by the Rodrigues formula Eq. (38), which requires to calculate the weight function $\rho(t)$ in (13). Then we end up with eqn. (12).
@@ IV Conclusion
The toy profile (2) proposed in LongMemory , which may or may not be similar to that of flyby encountered in nature, yields, for certain “magic” values of the wave parameter, a step-like, pure displacement behavior, consistently with the suggestions of Zel’dovich and Polnarev. The profile (2) can, moreover, be approximated by both the derived-Pöschl-Teller (3) and by the Scarf potentials (4). The latter supports analytic solutions, (11)-(12), constructed by the Nikiforov-Uvarov method and illustrated in FIG.1. The solutions are labeled by an integer $n$, cf. eqn (10).
The displacement effect (DM) arises when the wave zone contains $n$ half-waves as observed in DM-2 for the Gaussian and Pöschl-Teller profiles. The number of half-waves is equal to that of the zeros of the polynomial $y_{n}(t)$.
Our Scarf solution is similar to but slightly different from those found semi-analytically for derived-Pöschl-Teller DM-2 FIG.2. It is exact, while the one in DM-2 is only approximate, with error increasing for higher wave numbers.
@@
A recent paper MSUSY confirms our results using supersymmetric quantum mechanics.